OpenStudy (timaashorty):

Check my answer please. I think the answer is D. Is that correct or not? Identify the vertex and the axis of symmetry of the graph of the function y=3(x+2)^2-3. A. Vertex: (2,-3), Axis of symmetry: x=2 B. Vertex: (-2,-3), Axis of symmetry: x=-2 C. Vertex: (2,3), Axis of symmetry: x= 2 D. Vertex: (-2, 3), Axis of symmetry: x=-2

OpenStudy (jdoe0001):

\(\bf y=3(x+2)^2-3\implies y=3(x-(-2))^2+(-3)\)

OpenStudy (agent0smith):

Did you find the vertex? It will be (h, k), from @jdoe0001

OpenStudy (agent0smith):

From the pic he gave.

OpenStudy (timaashorty):

Thanx you guys (: It seems like the answer is A, but How do I tell about the axis of symmetry?

OpenStudy (agent0smith):

axis of symmetry is just the x-coordinate of the vertex. A isn't correct.

OpenStudy (timaashorty):

Jdoe, I don't understand the second comment, why did you turn the h, into -2?

OpenStudy (agent0smith):

Because the regular form is y = a(x - h)^2 + k and the vertex is (h, k) (notice the h changes sign) y = 3(x + 2)^2 - 3 so the vertex is (-2, -3)

OpenStudy (agent0smith):

And the x-coordinate of that vertex, is your axis of symmetry

OpenStudy (timaashorty):

So it's B ?

OpenStudy (agent0smith):

Yep :)

OpenStudy (timaashorty):

Thank you (:

OpenStudy (agent0smith):

welcome :)